OpenStudy (anonymous):

Step 1: |x − 2| + 3 = 7 Step 2: |x − 2| = 7 − 3 Step 3: |x − 2| = 4 Which of the following is a correct next step to solve the equation? x + 2 = −4 −x − 2 = 4 x + 2 = 4 x − 2 = −4

OpenStudy (astrophysics):

Well when you work with absolute values you can have two answers |x-2| = 4 can either give you x-2=4 or -(x-2)=4

OpenStudy (anonymous):

thank you sooo much!

OpenStudy (astrophysics):

Np